# NAMET: Robust Massive Model Editing via Noise-Aware Memory Optimization

### v1

@@ 6.4 Robustness against long prefixes (RQ3)
Finding 4. NAMET remains effective when edits are preceded by long prefixes. In Figure 6, NAMET consistently achieves the highest editing efficacy across all models. The maximum efficacy drop with 200-token prefixes is only 7% across the three models. In contrast, MEMIT’s efficacy decreases significantly from 84.75% to 66.50% on LLaMA2-7B, and from 94.2% to 82.25% on Deepseek-7B as prefix length increases. PMET also shows weak robustness, with efficacy drops of 17% and 30% on LLaMA2-7B and Deepseek-7B, respectively. Notably, all methods demonstrate strong robustness on Qwen2.5-7B, consistent with our earlier observation that Qwen2.5-7B is more suitable for robust batch editing. Nevertheless, NAMET exhibits the smallest efficacy drop (only 1.9%) when the prefix increases to 200 tokens, which is half that of the second-best method (MEMIT).
7 Limitations and Future Discussions
While NAMET demonstrates strong performance in massive knowledge editing, its potential in sequential knowledge editing remains under-explored. Specifically, recent work [6] on preserving model performance during sequential editing suggests that the hidden representations of pre-edited and post-edited models should remain similar. This contrasts with our findings, which show that more dispersed hidden representations enable the concurrent editing of a larger number of knowledge items. Future work should explore the balance between sequential and massive editing, as well as between uniform and dispersed hidden representations, in order to facilitate the application of model editing in real-world scenarios.
@@ 8 Conclusion
In this paper, we propose NAMET, a novel model editing method that enables stronger and more robust massive editing across various models and datasets. We first identify that the failures of existing methods in massive editing, particularly when evaluated with stricter metrics and context-rich settings, stem from conflicts between residual-key (r-k) and key-key (k-k) pairs. NAMET addresses this issue with a simple modification to MEMIT, introducing noise during the optimization of memory representations to promote more isolated r-k pairs, thereby reducing collisions. Extensive experiments on multiple base LLMs, including LLaMA2, LLaMA3, Deepseek, and Qwen, demonstrate that NAMET significantly outperforms existing methods in editing performance.
Appendix A Proof of Theorem 1
A.1 Problem Setting
We consider the following linear inverse problem that arises during massive editing in MLP layers. Suppose we aim to edit $N_{t}$ knowledge items to a pre-trained LLM. Each edit is associated with:
  • A key vector $k_{i}\in\mathbb{R}^{n}$, which encodes the the factual subject;
  • A residual vector $r_{i}\in\mathbb{R}^{m}$, representing the target object.
Let:
$S_{i}:=k_{i}k_{i}^{T}\in\mathbb{R}^{n\times n},\quad T_{i}:=r_{i}k_{i}^{T}\in\mathbb{R}^{m\times n}.$ (12)
That is, each $S_{i}$ is the outer product of the key with itself (key-key embedding), and $T_{i}$ is the outer product of residual and key (residual-key embedding). The update on MLP layers aims to solve the following equation:
$\Delta\left(C_{p}+\sum_{i=1}^{N_{t}}S_{i}\right)=\sum_{i=1}^{N_{t}}T_{i},$ (13)
where $C_{p}$ denotes the preserved key embeddings that should remain unaffected by the update. This leads to the additional constraint $\Delta C_{p}=0$, which transforms Equation 13 into:
$\Delta S=T,\quad\text{where}\quad S:=\sum_{i=1}^{N_{t}}S_{i},\quad T:=\sum_{i=1}^{N_{t}}T_{i}.$ (14)
We consider the minimum-norm solution for the update:
$\Delta:=TS^{+},$ (15)
where $S^{+}$ denotes the Moore-Penrose pseudoinverse of $S$. The reconstructed target for item $i$ is then:
$\hat{T}_{i}:=\Delta S_{i},$ (16)
and the reconstruction loss is then:
$RL_{i}:=\hat{T}_{i}-T_{i}=\Delta S_{i}-T_{i}.$ (17)
A.2 Error Decomposition
We begin by expanding:
$\Delta=\left(\sum_{j}T_{j}\right)S^{+}\quad\Rightarrow\quad RL_{i}=\left(\sum_{j}T_{j}\right)S^{+}S_{i}-T_{i}.$ (18)
This can be decomposed as:
$RL_{i}=(T_{i}S^{+}S_{i}-T_{i})+\sum_{j\neq i}T_{j}S^{+}S_{i},$ (19)
where:
  • The first term, $T_{i}S^{+}S_{i}-T_{i}$, is the self-alignment bias, reflecting how well $T_{i}$ is reconstructed through its own $S_{i}$;
  • The second term, $\sum_{j\neq i}T_{j}S^{+}S_{i}$, is the interference error, reflecting how other items’ r-k embeddings $T_{j}$ leak into $\hat{T}_{i}$ due to overlap in k-k embeddings.
A.3 Impact of Non-Orthogonality among $\{S_{i}\}$ (Key-Key Collisions)
Assume $S$ is full-rank and well-conditioned, so we approximate:
$S^{+}\approx S^{T}(SS^{T})^{+}.$ (20)
Then:
$T_{j}S^{+}S_{i}\approx T_{j}S^{T}(SS^{T})^{+}S_{i}.$ (21)
This term is large when $S_{i}$ shares directions with other $S_{j}$ (i.e., $S_{i}$ is not orthogonal to $S_{j}$). More precisely, we can express:
$S^{+}S_{i}=\sum_{k}w_{ik}S_{k}^{T},\quad\Rightarrow\quad T_{j}S^{+}S_{i}=\sum_{k}w_{ik}T_{j}S_{k}^{T}.$ (22)
Thus, if $S_{i}$ is not orthogonal to other $S_{k}$, then $S_{i}$ shares eigen-directions with $S_{k}$, and $\hat{T}_{i}$ becomes vulnerable to contamination from $T_{j}$.
A.4 Impact of Non-Orthogonality among $\{T_{i}\}$ (Residual-Key Collisions)
Now suppose r-k embeddings are redundant across edits:
$T_{i}=\bar{T}_{i}+\sum_{j\neq i}\gamma_{ij}T_{j}.$ (23)
where $\bar{T}_{i}$ reflects the real r-k embedding of target item $i$. Then:
$T=\bar{T}_{i}+\sum_{j\neq i}(1+\gamma_{ij})T_{j},\quad\Rightarrow\quad\Delta=TS^{+}=\bar{T}_{i}S^{+}+\sum_{j\neq i}(1+\gamma_{ij})T_{j}S^{+}.$ (24)
Hence,
$\hat{T}_{i}=\Delta S_{i}=\bar{T}_{i}S^{+}S_{i}+\sum_{j\neq i}(1+\gamma_{ij})T_{j}S^{+}S_{i}.$ (25)
The interference term is scaled both by k-k overlap (through $S^{+}S_{i}$) and by r-k redundancy (via $\gamma_{ij}$).
A.5 Conclusion: Double Embedding Collision Amplifies Error
We restate the main result of this section as Theorem 1.
Theorem 1. Given keys $k_{i}\in\mathbb{R}^{n}$ and residuals $r_{i}\in\mathbb{R}^{m}$, define $S_{i}=k_{i}k_{i}^{T}$ and $T_{i}=r_{i}k_{i}^{T}$. Let $S=\sum_{i}S_{i}$, $T=\sum_{i}T_{i}$, and $\Delta=TS^{+}$. Then the reconstruction error $RL_{i}:=\Delta S_{i}-T_{i}$ can be decomposed as:
$RL_{i}=(T_{i}S^{+}S_{i}-T_{i})+\sum_{j\neq i}T_{j}S^{+}S_{i}.$ (26)
The magnitude of $RL_{i}$ is correlated with both:
$\sum_{j\neq i}\cos(k_{i}k_{i}^{T},k_{j}k_{j}^{T})\quad\text{and}\quad\sum_{j\neq i}\cos(r_{i}k_{i}^{T},r_{j}k_{j}^{T}),$ (27)
which we refer to as the embedding collision scores in the key-key and residual-key embeddings, respectively. These collisions hinder the disentanglement of individual edits in the shared parameter space, thereby degrading editing performance under massive updates.
Appendix B Detailed Experiment Setup
@@ B.1 Datasets and Baselines
  • CounterFact [19] is a benchmark for evaluating factual knowledge localization and editing in LLMs. It contains 21,917 entries that describe the named entities along with their counterfactual variations. Model editing techniques could be evaluated in terms of editing efficacy, generalization, and locality. The benchmark also contains generation prompts to test the model’s generation ability after editing.
  • ZsRE [14] is a question-answering (QA) benchmark designed to evaluate zero-shot relation extraction capabilities of language models. Entries in the benchmark consist of a subject entity along with an answer as the editing target. The benchmark also includes paraphrased questions for testing generalization ability and irrelevant questions for evaluating the locality of editing techniques.
  • Wiki-recent [37] contains 1,266 entries of triplets that have been added into WikiData after July 2022. The benchmark enables insertion for models that were trained prior to the introduction of these facts. This simulates the cases of editing outdated models with newly introduced facts. Model editing techniques are evaluated in terms of editing efficacy, portability, and locality. Here, portability emphasizes whether the edited model could reason about the downstream effects of facts when they are inserted into the model.
@@
  • MEND [21] requires extra parameters for efficiently editing pretrained LLMs. It introduces a set of small auxiliary networks that transform standard fine-tuning gradients into low-rank updates, enabling fast and localized edits without retraining the entire model. This approach offers a scalable solution for post-hoc model editing, avoiding the overfitting issue of traditional fine-tuning methods.
  • ROME [19] performs factual knowledge editing by directly modifying the feed-forward weights in specific layers of LLMs. It first identifies that factual knowledge is primarily stored in mid-layer feed-forward modules, thereby demonstrating the feasibility of editing model parameters to update internal knowledge. The method then updates these weights to encode specific factual associations. ROME achieves precise insertion of new facts with minimal interference to unrelated knowledge.
  • MEMIT [20] is designed to efficiently update LLMs with thousands of factual associations simultaneously. Building upon ROME, MEMIT employs a least-squares optimization over multiple key-value memory components, ensuring high specificity and minimal interference with unrelated knowledge. It further distributes the updates across multiple layers, which helps reduce the impact on the model’s general capabilities.
  • PMET [15] is a method designed to enhance the precision of knowledge updates in large language models. Unlike prior approaches that treat transformer layer (TL) hidden states as direct inputs of the feed-forward network (FFN), PMET recognizes that these hidden states also encompass information from multi-head self-attention (MHSA) and residual connections. PMET proceeds to simultaneously optimize MHSA and FFN hidden states and use the optimized TC hidden states of FFN to precisely update FFN weights. This approach enables more accurate and efficient model editing, preserving the integrity of the model’s existing knowledge while incorporating new information.
  • ALPHAEDIT [6] preserves knowledge in LLMs during sequential updates by projecting updates onto the null space of the preserved knowledge. This could ensure that new modifications do not interfere with previously stored information. This approach maintains the integrity of the model’s existing knowledge while enabling precise edits.
B.2 Metrics

### v4

@@ 6.4 Robustness against long prefixes (RQ3)
Finding 5. EAMET Remains Effective When Edits Are Preceded by Long Prefixes. In Figure˜5, EAMET achieves the highest editing efficacy across all models, with at most a 7% drop at 200-token prefixes. In contrast, MEMIT suffers a much larger decline, from 84.75% to 66.50% on LLaMA2-7B and from 94.2% to 82.25% on DeepSeek-7B. Notably, all methods demonstrate strong robustness on Qwen2.5-7B, consistent with our earlier observation that Qwen2.5-7B is more suitable for robust batch editing. Nevertheless, EAMET exhibits the smallest efficacy drop (only 1.9%) when the prefix increases to 200 tokens, which is half that of the second-best method (MEMIT).
6.5 Robustness under Multiple Edits of the Same Subject (RQ4)
Figure 6: Editing performance of different methods across varying numbers of facts per subject.
We evaluate the robustness of editing methods when multiple facts concerning the same subject are edited simultaneously. Specifically, we simultaneously edit 10,000 facts from ZsRE dataset, and only evaluate samples whose subject is associated with multiple facts. We group subjects according to the number of associated samples and examine how rewrite accuracy varies with this number. Experiments are conducted on LLaMA2-7B, DeepSeek-7B, and Qwen2.5-7B.
Finding 6. EAMET Remains Effective When Multiple Facts of the Same Subject Are Edited Simultaneously. Figure˜6 shows that EAMET consistently achieves the highest editing efficacy across nearly all settings. Its performance remains stable when editing multiple samples associated with the same subject. In contrast, other methods exhibit a clear decline in efficacy as the number of facts per subject increases, which ultimately results in degraded performance on the overall massive editing task.
@@ 7 Conclusion
In this paper, we propose EAMET, a novel model editing method that enables stronger and more robust massive editing across various models and datasets. We first identify that the failures of existing methods in both effectiveness and robustness of massive editing stem from misalignment between the space of key and residual embeddings. EAMET addresses this issue by progressively aligning the key and residual embedding space when optimizing target memory for each fact. The aligned embeddings increase both the capacity and robustness of massive editing. Extensive experiments on multiple base LLMs, including LLaMA2, LLaMA3, Deepseek, and Qwen, demonstrate that EAMET significantly outperforms existing methods in editing performance and robustness.
Acknowledgments
The HKUST authors were supported in part by a RGC GRF grant under the contract 16214723, research fund provided by CMHK and ZTE, and a HKUST Bridge The Gap fund BGF.001.2025. We are grateful to the anonymous reviewers for their valuable comments. We thank HKUST Fok Ying Tung Research Institute and National Supercomputing Center in Guangzhou Nansha Sub-center for computational resources.
Ethics Statement
This work does not raise any specific ethical concerns. Its primary goal is to advance research on model editing and to offer a new perspective on this problem.
Reproducibility Statement
All experiments in this paper are reproducible. The code and datasets are publicly available at the GitHub repository: https://github.com/ybdai7/EAMET-massive-editing.
Appendix A LLM Usage Statement
We used LLMs solely to assist in drafting and polishing the writing of this paper, without any other purposes.
Appendix B Proof of Theorem 1
Notation Setup. Let $K=[k_{1},\dots,k_{N}]\in\mathbb{R}^{d\times N}$ and $R=[r_{1},\dots,r_{N}]\in\mathbb{R}^{d\times N}$, and define
$M:=\sum_{i=1}^{N}k_{i}k_{i}^{\top}=KK^{\top},\qquad N:=\sum_{i=1}^{N}r_{i}k_{i}^{\top}=RK^{\top}.$ (16)
Assume that $M$ is (pseudo-)invertible and define the closed-form solution $\Delta=NM^{+}=RK^{\top}(KK^{\top})^{+}$, so that $\Delta\sum_{i}k_{i}k_{i}^{\top}=\sum_{i}r_{i}k_{i}^{\top}$.
Derivation of the Column Expansion $\Delta k_{i}=\sum_{j}\beta_{ij}r_{j}$. By definition,
$\Delta=\sum_{j=1}^{N}r_{j}k_{j}^{\top}M^{+}=RK^{\top}M^{+}.$ (17)
Applying $\Delta$ to a column $k_{i}$ gives
$\Delta k_{i}=\sum_{j=1}^{N}r_{j}k_{j}^{\top}M^{+}k_{i}.$ (18)
Setting $\beta_{ij}:=k_{j}^{\top}M^{+}k_{i}$, we immediately obtain
$\Delta k_{i}=\sum_{j=1}^{N}\beta_{ij}r_{j}.$ (19)
This formula provides an explicit linear combination of residual embeddings $r_{j}$ that reconstructs $\Delta k_{i}$, with coefficients $\beta_{ij}$ determined by the key embeddings and the pseudo-inverse of $M$.
Reconstruction Residual and Neighborhood Decomposition. For each knowledge item $i$, define the reconstruction residual $\mathrm{e}_{i}:=\Delta k_{i}-r_{i}$. Suppose that $r_{i}$ can be approximately reconstructed from its neighbors with nonnegative weights $q_{ij}$ for $j\neq i$, i.e.,
$r_{i}=\sum_{j\neq i}q_{ij}\,r_{j}+\varepsilon_{i},\qquad q_{ij}\geq 0,\;\sum_{j\neq i}q_{ij}=1,$ (20)
where $\varepsilon_{i}$ denotes the residual error. Substituting this decomposition into $\mathrm{e}_{i}$ gives
$\mathrm{e}_{i}=\sum_{j\neq i}(\beta_{ij}-q_{ij})\,r_{j}+\beta_{ii}r_{i}-\varepsilon_{i}.$ (21)
Bounding the Reconstruction Residual. Taking norms and applying the triangle inequality yields
$\|\mathrm{e}_{i}\|\leq\sum_{j\neq i}|\beta_{ij}-q_{ij}|\,\|r_{j}\|+|\beta_{ii}|\,\|r_{i}\|+\|\varepsilon_{i}\|.$ (22)
To relate the first term to embedding alignment, we construct a probability vector $p_{i}$ from the positive parts of the coefficients $\beta_{ij}$ (for $j\neq i$):
$s_{ij}:=\max\{\beta_{ij},0\},\qquad S_{i}:=\sum_{j\neq i}s_{ij},\qquad p_{ij}:=\frac{s_{ij}}{S_{i}}.$ (23)
Defining $C_{i}:=\sum_{j\neq i}\|r_{j}\|$, one can show that
$\sum_{j\neq i}|\beta_{ij}-q_{ij}|\,\|r_{j}\|\leq C_{i}\,\mathrm{TV}(p_{i},q_{i}),$ (24)
up to negligible contributions from negative $\beta_{ij}$ that can be absorbed into $C_{i}$. Here, $\mathrm{TV}(p_{i},q_{i})$ is the total variation (TV) distance between two discrete distributions $p_{i}$ and $q_{i}$:
$\mathrm{TV}(p,q):=\tfrac{1}{2}\sum_{j}|p_{j}-q_{j}|.$ (25)
Finally, applying Pinsker’s inequality $\mathrm{TV}(p_{i},q_{i})\leq\sqrt{\tfrac{1}{2}\,\mathrm{KL}(q_{i}\|p_{i})}$ (Cover and Thomas, 1999) gives
$\|\mathrm{e}_{i}\|\leq C_{i}\,\sqrt{\frac{1}{2}\,\mathrm{KL}(q_{i}\|p_{i})}+|\beta_{ii}|\,\|r_{i}\|+\|\varepsilon_{i}\|.$ (26)
Identifying $q_{i}=P_{r}^{(i)}$ and $p_{i}=P_{k}^{(i)}$ with the kernel-normalized neighborhood distributions from Definition 1 yields the embedding-alignment bound stated in the main text:
$\|\mathrm{e}_{i}\|\leq C_{i}\sqrt{\frac{1}{2}\,\mathcal{A}(i)}+|\beta_{ii}|\,\|r_{i}\|+\|\varepsilon_{i}\|.$ (27)
If $M$ is singular, replace $M^{+}$ with the Moore–Penrose pseudoinverse. The contributions from negative $\beta_{ij}$ or scaling factors can usually be absorbed into $C_{i}$. In the ideal case of perfect neighborhood alignment $\mathcal{A}(i)=0$, negligible self-weight $\beta_{ii}=0$, and vanishing residual $\varepsilon_{i}=0$, we recover $\mathrm{e}_{i}=0$.
Appendix C Detailed Analysis of Corollary 2
We provide a detailed theoretical analysis that develops Corollary 2.
Residual Decomposition. Recall that the reconstruction residual can be written as
$e_{i}=\Delta k_{i}-r_{i}=\sum_{j=1}^{N}\beta_{ij}r_{j}-r_{i},\qquad\beta_{ij}:=k_{j}^{\top}\!\Big(\sum_{\ell=1}^{N}k_{\ell}k_{\ell}^{\top}\Big)^{\!-1}\!k_{i}.$ (28)
Partition the index set into the subject cluster $\mathcal{S}$ (samples sharing the subject with $i$) and the remainder $\mathcal{T}$. Then
$\Delta k_{i}=\beta_{ii}r_{i}+\sum_{j\in\mathcal{S},j\neq i}\beta_{ij}r_{j}+\sum_{j\in\mathcal{T}}\beta_{ij}r_{j}.$ (29)
Empirically and theoretically, only coefficients $\beta_{ij}$ for $j\in\mathcal{S}$ become significant, while cross-subject coefficients remain negligible since embeddings from different subjects are nearly orthogonal in key space and thus contribute little to the reconstruction. Hence reconstruction is dominated by the same-subject neighborhood.
Effect of Adding Same-Subject Samples. Let $K=\sum_{\ell}k_{\ell}k_{\ell}^{\top}$ denote the Gram matrix. Suppose we add one additional key $k_{j}$ (with $j\in\mathcal{S},j\neq i$). By the Woodbury identity(Horn and Johnson, 1985),
$(K+k_{j}k_{j}^{\top})^{-1}=K^{-1}-\frac{K^{-1}k_{j}k_{j}^{\top}K^{-1}}{1+k_{j}^{\top}K^{-1}k_{j}}.$ (30)
Consequently, the updated self-weight becomes
$\beta_{ii}^{\text{new}}=k_{i}^{\top}(K+k_{j}k_{j}^{\top})^{-1}k_{i}=\beta_{ii}-\frac{(k_{i}^{\top}K^{-1}k_{j})^{2}}{1+k_{j}^{\top}K^{-1}k_{j}}.$ (31)
Thus $\beta_{ii}$ monotonically decreases as more same-subject vectors are included. The lost weight is redistributed into off-diagonal terms $\beta_{ij}$, consistent with our empirical observation in Figure 3.
Connection to Alignment. Using the decomposition in equation 29, the residual can be expressed as
$e_{i}=\sum_{j\in\mathcal{S}}\beta_{ij}(r_{j}-r_{i})+\sum_{j\in\mathcal{T}}\beta_{ij}r_{j}.$ (32)
Applying the triangle inequality yields
$\|e_{i}\|\leq\sum_{j\in\mathcal{S}}|\beta_{ij}|\,\|r_{j}-r_{i}\|+\sum_{j\in\mathcal{T}}|\beta_{ij}|\,\|r_{j}\|.$ (33)
The first term depends on the dispersion of responses within the same subject. This dispersion is controlled by the alignment measure $\mathcal{A}(i)$: when $\mathcal{A}(i)$ is small, the responses $\{r_{j}:j\in\mathcal{S}\}$ are tightly clustered around $r_{i}$, so even a redistribution of weight from $\beta_{ii}$ to other $\beta_{ij}$ produces only minor error. Conversely, when $\mathcal{A}(i)$ is large, intra-subject responses differ substantially, and the redistributed weights amplify reconstruction error.
Combining equation 31 and equation 33, we conclude that co-editing additional same-subject samples (i) monotonically decreases $\beta_{ii}$, (ii) redistributes weight into off-diagonal $\beta_{ij}$, and (iii) yields residuals bounded by the intra-subject alignment $\mathcal{A}(i)$. Therefore, massive editing performance crucially depends on the degree of alignment within the subject cluster.
Appendix D Detailed Experiment Setup
@@ D.1 Datasets and Baselines
  • CounterFact (Meng et al., 2022) is a benchmark for evaluating factual knowledge localization and editing in LLMs. It contains 21,917 entries that describe the named entities along with their counterfactual variations. Model editing techniques could be evaluated in terms of editing efficacy, generalization, and locality. The benchmark also contains generation prompts to test the model’s generation ability after editing.
  • ZsRE (Levy et al., 2017) is a question-answering (QA) benchmark designed to evaluate zero-shot relation extraction capabilities of language models. Entries in the benchmark consist of a subject entity along with an answer as the editing target. The benchmark also includes paraphrased questions for testing generalization ability and irrelevant questions for evaluating the locality of editing techniques.
  • Wiki-recent (Zhang et al., 2024) contains 1,266 entries of triplets that have been added into WikiData after July 2022. The benchmark enables insertion for models that were trained prior to the introduction of these facts. This simulates the cases of editing outdated models with newly introduced facts. Model editing techniques are evaluated in terms of editing efficacy, portability, and locality. Here, portability emphasizes whether the edited model could reason about the downstream effects of facts when they are inserted into the model.
@@
  • MEND (Mitchell et al., 2022a) requires extra parameters for efficiently editing pretrained LLMs. It introduces a set of small auxiliary networks that transform standard fine-tuning gradients into low-rank updates, enabling fast and localized edits without retraining the entire model. This approach offers a scalable solution for post-hoc model editing, avoiding the overfitting issue of traditional fine-tuning methods.
  • ROME (Meng et al., 2022) performs factual knowledge editing by directly modifying the feed-forward weights in specific layers of LLMs. It first identifies that factual knowledge is primarily stored in mid-layer feed-forward modules, thereby demonstrating the feasibility of editing model parameters to update internal knowledge. The method then updates these weights to encode specific factual associations. ROME achieves precise insertion of new facts with minimal interference to unrelated knowledge. When evaluating using ROME, we edit all facts sequentially with batch size 1, as it does not support batch editing.
  • MEMIT (Meng et al., 2023) is designed to efficiently update LLMs with thousands of factual associations simultaneously. Building upon ROME, MEMIT employs a least-squares optimization over multiple key-value memory components, ensuring high specificity and minimal interference with unrelated knowledge. It further distributes the updates across multiple layers, which helps reduce the impact on the model’s general capabilities.
  • PMET (Li et al., 2025) is a method designed to enhance the precision of knowledge updates in large language models. Unlike prior approaches that treat transformer layer (TL) hidden states as direct inputs of the feed-forward network (FFN), PMET recognizes that these hidden states also encompass information from multi-head self-attention (MHSA) and residual connections. PMET proceeds to simultaneously optimize MHSA and FFN hidden states and use the optimized TC hidden states of FFN to precisely update FFN weights. This approach enables more accurate and efficient model editing, preserving the integrity of the model’s existing knowledge while incorporating new information.
  • ALPHAEDIT (Fang et al., 2024) preserves knowledge in LLMs during sequential updates by projecting updates onto the null space of the preserved knowledge. This could ensure that new modifications do not interfere with previously stored information. This approach maintains the integrity of the model’s existing knowledge while enabling precise edits.
D.2 Metrics